OpenStudy (anonymous):

how to solve d(xy) integrals? for example d(xy)/(x^2+y^2)

OpenStudy (freckles):

is that real the question?

OpenStudy (freckles):

really*

OpenStudy (freckles):

you want to integrate d(xy)/(x^2+y^2)?

OpenStudy (anonymous):

yes i was solivng diferential equation and got this

ganeshie8 (ganeshie8):

whats the differential equation ?

OpenStudy (p0sitr0n):

you take one variable as constant and integrate with respect to the other one

OpenStudy (p0sitr0n):

doesnt matter which one first, theyre interchangeable

ganeshie8 (ganeshie8):

are these partials ?

OpenStudy (p0sitr0n):

assuming d(xy)= dxdy

OpenStudy (primeralph):

No.

OpenStudy (anonymous):

dxy is definitely not dxdy and about equation it will take a bit to find which 1 it was but anyway is there any way to solve such an equation?

ganeshie8 (ganeshie8):

could you post the original differential equation ?

OpenStudy (anonymous):

its an equation where i tried to pick an integral factor (x+x^2+y^2)dy - ydx=0 and got that the factor is e^ (-2d(xy)/(xx+yy))

ganeshie8 (ganeshie8):

btw, you cannot take integral of an unknown function "y" with respect to another variable "x"

ganeshie8 (ganeshie8):

is that linear ?

ganeshie8 (ganeshie8):

nope. so integrating factor thing may not work right ?

OpenStudy (anonymous):

well if u have 1d(xy) the answer is xy+c so i guess there should be some way to solve harder integrals with dxy

OpenStudy (anonymous):

theres task in book to find a factor for this equation

ganeshie8 (ganeshie8):

not sure how to proceed.. wolfram gives a simple solution http://www.wolframalpha.com/input/?i=%28x%2Bx%5E2%2By%5E2%29dy+-+ydx%3D0

OpenStudy (anonymous):

well anyway i was just wondering if such an integral is solvable and even real anywhere else

ganeshie8 (ganeshie8):

@eliassaab

OpenStudy (anonymous):

ok thanks

ganeshie8 (ganeshie8):

@wio speaks this language of d(xy)'s and stuff.. . :)

ganeshie8 (ganeshie8):

what is \(\large \int \frac{d(y/x)} {1 + (y/x)^2}\) ?

OpenStudy (anonymous):

um?

ganeshie8 (ganeshie8):

lol in which subject u get to do these crazy stuff ?

OpenStudy (anonymous):

i dont think if you can connect mine integral to this 1 but this 1 should be arctan the subject is differential equations

ganeshie8 (ganeshie8):

mhmm

OpenStudy (anonymous):

\[ d(xy) = (xy)'dx = (y+xy')dx \]

OpenStudy (anonymous):

But that won't help you much because you'll have a rogue \(y\) term. So consider a transformation.

OpenStudy (anonymous):

why is its (xy)'dx and not dy or whatever else? is it because y(x)?

OpenStudy (anonymous):

I didn't say you couldn't do \[ d(xy) = (xy)'dy = (x'y+x)dy \]

OpenStudy (anonymous):

in this case \(x' = dx/dy\)

OpenStudy (anonymous):

ohhh right

OpenStudy (anonymous):

well thanks now i know how to deal with these integrals :D

OpenStudy (anonymous):

I think you'll still have to use a transformation though.

OpenStudy (anonymous):

yeah but the equation is of a lesser issue here as i just was wondering how to deal with these integrals because i have never had need to solve them

OpenStudy (anonymous):

You mean actually doing the integral is less of a problem?

OpenStudy (anonymous):

nope just in this case was looking for how to solve the integral

OpenStudy (anonymous):

by the way is there any problem that we have also \[\frac{ 1 }{ x^2+y^2 }\] next to dxy